# Crowd-SAM: SAM as a Smart Annotator for Object Detection in Crowded Scenes

### v1

@@ 4.3 Ablation Studies
Ablation on the modules. We conduct ablation studies on the key components of Crowd-SAM - foreground location, EPS, and PWD-Net - to validate their effectiveness. In Tab. 6, the performance on AP significantly drops by 7.4% and recall by 8.5% when the FG location is removed, indicating the importance of restricting foreground areas. As for EPS, once it is removed, the AP drops by 0.6%. We suppose EPS not only accelerates the sampling process of dense prompts but also helps focus on the difficult part of the images, which has ambiguous semantics. We also compare EPS with other batch iterators in  Tab. 7. Moreover, we find the PWD-Net is indispensable and when removed, the AP dramatically drops to only 17.0%. Finally, we point out that multi-cropping is a strong method to enhance the performance which contributes 7.4 % AP to the final performance. Overall, these results prove that all components are essential.
@@
Table 7: Comparisons of different sampler. Full means using all prompts. OOM represents out-of-memory error which occurs when the GPU memory is all consumed.
Ablation on EPS. We conduct a comprehensive study on EPS by comparing it with several variants. We forward each image only once to avoid the side effects caused by multi-crop. We use a default sampler that iterates through all training samples and a random iterator with a halting threshold $K$ as counter-parts. As reported in  Tab. 7, AP and Recall increase with the grid size for all three samplers. However, the default sampler meets an out-of-memory error when the grid size reaches 128, preventing it from being adopted in this setting. As for the random sampler, its performance is constrained by K and a larger grid size than 64 only leads to limited improvement, e.g. 0.1% AP. On the contrary, our EPS can benefit from a much larger grid size and achieves a better performance especially when the grid size is large.
@@
Table 8: Ablation studies on the design of PWD-Net. ${\cal M}$, ${\cal U}$ and ${\cal O}$ represent mask token, IoU token, and semantic token, respectively. In the column of the IoU head, $F$ means to freeze the original IoU head and train a parallel one and $T$ means tuning the original IoU head.
Ablation on PWD-Net. We compare PWD-Net with variants that replace some tokens with a full-zero placeholder. We also compare with two designs, tuning the IoU head directly or learning a parallel IoU head. As shown in  Tab. 8, all three tokens, i.e. Mask Token ${\cal M}$, IoU Token ${\cal U}$, and Semantic Token ${\cal O}$, contribute to the final result. Particularly, once ${\cal M}$ is removed, the AP drops by 40.0%, which is a catastrophic decline. This degradation suggests that the Mask Token contains the essential shape-aware feature that is essential for the part-whole discrimination task. Notably, the AP drops by 2.8% when we tune the pre-trained IoU Head, suggesting that it might overfit the few labeled images. We believe by freezing the SAM’s IoU head, the PWD-Net can benefit more from the shape-aware knowledge learned on massive segmentation data.
Figure 4: Qualitative comparisons between our proposed Crowd-SAM (a) and De-FRCN (b). We also plot the generated masks and GT boxes (c) and our prompts filtering results (d), where red points are kept and gray points are filtered.
@@ 5 Conclusion
This paper introduces Crowd-SAM, a SAM-based framework for object detection and segmentation in crowded scenes, designed to streamline the annotation process. For each image, Crowd-SAM generates dense prompts for high recall and uses EPS to prune redundant prompts. To achieve accurate detection in occlusion cases, Crowd-SAM employs PWD-Net that leverages several informative tokens to select the best-fitting mask. Combined with the proposed methods, Crowd-SAM achieves 78.4 AP on Crowd Human, comparable to full-supervised detectors, validating that object detection in crowded scenes can benefit from foundation models like SAM in data efficiency.

### v2

@@ 4.3 Ablation Studies
Ablation on Modules. We conduct ablation studies on the key components of Crowd-SAM, including foreground location, EPS, and PWD-Net, to validate their effectiveness. In Tab. 6, the performance in AP significantly drops by 7.4% and recall by 8.5% when FG location is removed, indicating the importance of restricting foreground areas. As for EPS, once it is removed, the AP drops by 0.6%. We suppose that EPS not only accelerates the sampling process of dense prompts but also helps focus on the difficult part of the image, which conveys ambiguous semantics. We also compare EPS with some other batch iterators in Tab. 7. Besides, we find that PWD-Net is indispensable and when removed, the AP dramatically drops to 17.0%. Finally, we point out that multi-cropping is a strong trick to enhance the performance which contributes 7.0% AP to the final performance. Overall, these results prove that all the components are essential.
Table 7: Comparison (%) of different samplers on CrowdHuman [37] val. Full means using all prompts. OOM represents out-of-memory errors which occur when the GPU memory is all consumed.
@@
Ablation on EPS. We conduct a comprehensive study on EPS by comparing it with several variants. We forward each image only once to avoid the extra latency caused by multi-crop. We use a default sampler that iterates through all training samples and a random iterator with a halting threshold $K$ as counter-parts. As reported in Tab. 7, the AP and Recall increase with the grid size for all three samplers. However, the default sampler suffers an out-of-memory error when the grid size reaches 128, preventing it from being adopted in this setting. As for the random sampler, its performance is constrained by $K$ and a grid size larger than 64 only leads to limited improvement, e.g. 0.1% AP. On the contrary, our EPS benefits from a much larger grid size and achieves a better result.
Table 8: Ablation results (%) on the design of PWD-Net. ${\cal M}$, ${\cal U}$, and ${\cal O}$ represent the mask token, IoU token, and semantic token, respectively. For the IoU head, $F$ means freezing the original IoU head and training a parallel one, and $T$ indicates tuning the original IoU head.
@@
Ablation on PWD-Net. We compare PWD-Net to the variants that replace some tokens with a full-zero placeholder. We also consider two designs, directly tuning the IoU head or learning a parallel IoU head. As shown in  Tab. 8, all the three tokens, i.e. Mask Token ${\cal M}$, IoU Token ${\cal U}$, and Semantic Token ${\cal O}$, contribute to the final result. Particularly, once ${\cal M}$ is removed, the AP drops by 40.0%, which is a catastrophic decline. This degradation suggests that the mask token contains the shape-aware feature that is essential for the part-whole discrimination task. Notably, the AP drops by 2.8% when we tune the pre-trained IoU Head, suggesting that it is prone to overfit the few labeled images. By freezing the IoU head of SAM, PWD-Net can benefit more from the shape-aware knowledge learned from massive segmentation data.
Figure 4: Qualitative comparison between Crowd-SAM (a) and De-FRCN (b). Crowd-SAM predictions are more accurate especially in the boundaries of persons. We also plot the GT boxes (blue rectangles) and the generated masks (yellow regions), which are of high quality (c). In (d), we plot our prompt filtering results, where preserved prompts (red points) are much fewer than the removed ones (gray points). Zoom in for a better view.
@@ 5 Conclusion
This paper proposes Crowd-SAM, a SAM-based framework, for object detection and segmentation in crowded scenes, designed to streamline the annotation process. For each image, Crowd-SAM generates dense prompts for high recall and uses EPS to prune redundant prompts. To achieve accurate detection in occlusion cases, Crowd-SAM employs PWD-Net which leverages several informative tokens to select the masks that best fit. Combined with the proposed modules, Crowd-SAM achieves 78.4% AP on CrowdHuman, comparable to full-supervised detectors, validating that object detection in crowded scenes can benefit from foundation models like SAM with data efficiency.